OpenStudy (anonymous):

Telephone bill decreased from about $50 a month to about $47. What is the percent of decrease?

OpenStudy (anonymous):

6%

OpenStudy (anonymous):

actual decrease is $3 so take \(\frac{3}{50}\) and turn it in to a percent

OpenStudy (anonymous):

easy enough in this case because \(\frac{3}{50}=\frac{6}{100}=6\%\)

OpenStudy (anonymous):

Thanks @satellite73 :*

OpenStudy (anonymous):

yw